OpenStudy (anonymous):

Find the Laplace transform of (1-e^-t)/t

OpenStudy (anonymous):

From Mathematica:\[\text{LaplaceTransform}\left[\frac{1-e^{-t}}{t},t,s\right]\to \text{Log}\left[1+\frac{1}{s}\right] \] Never had a course in the subject. Sorry.

OpenStudy (anonymous):

i know the answer too,.. i just needed to know how to do it.. it's okay. thanks for your help!